OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 5x + 4 = 0 Part A: Describe the solution(s) to the equation by just determining radicand. Show your work.

OpenStudy (amistre64):

so its asking you to consider the quadratic formula, but just the part that is the square root.

OpenStudy (anonymous):

so just break it down to (x-?) (x+?)=0? @amistre64

OpenStudy (anonymous):

or it might be ++ instead of +-

OpenStudy (amistre64):

its not asking for the solutions, but rather what kind of solutions we would expect given that the radicand (the discriminant) is positive, negative, or 0

OpenStudy (amistre64):

recall the quad formula, given: ax^2 + bx + c = 0 \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] now if: b^2 - 4ac is positive, negative, or zero tells us the type of roots we will expect to get

OpenStudy (amistre64):

so, 5^2 -4(1)(4) = 25-16 its positive, so what type of roots do we expect? 2 real and different 2 real and the same or a complex, conjugate pair

OpenStudy (anonymous):

2real and diffrent is what we will expect i think

OpenStudy (amistre64):

thats correct

OpenStudy (amistre64):

if its 0, then vertex + 0 is the same as vertex - 0 if its negative, then well, sqrt(-) has no real solutions

OpenStudy (anonymous):

okay now u confused me

OpenStudy (haseeb96):